OpenStudy (anonymous):

Solve. 4 + x/7 = 2 A) -14 B) 10 Eliminate C) 12 D) 42

sammixboo (sammixboo):

First subtract 4 on both sides of the equation 4 + -4 x/7 = 2 - 4 x/7 = 2 - 4 Can you tell me what is 2 - 4?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Solve. 4 + x/7 = 2 Subtract 4 from both sides x/7 = -2 Then multiply both sides by 7 x = -14 Answer is -14

OpenStudy (anonymous):

Thanks! both of you

sammixboo (sammixboo):

No problem!

OpenStudy (anonymous):

You're just solving for the variable. All you have to do is eliminate all other things surround the variable until it stands alone on that side of the equation.

OpenStudy (anonymous):

So it will look like x = 5 or x = 11

OpenStudy (anonymous):

Oh ok :)